OpenStudy (anonymous):

how do i solve a first order DE using Eulers method? how do i get the software first?? help...

OpenStudy (anonymous):

I'm confused, u want to do manually or u want to do with software?

OpenStudy (anonymous):

this is a numerical method for equations of form dy/dx = f(x,y) given an initial condition y = ya eulers method generates a sequence of approximations yn =approx y(xn) and y n+1 is obtained from yn by the formula y n+1 = yn + hf(xn, yn ), n = 1,2,3,..... where y1 = ya and xn+1 = xn + h and h is a positive step-size for example y' = (1-x)y + cos x, y(0) = 1 eulers method forms y1 = 1 + 2h as an approximation to y(h)